# VHE gamma-ray intranight variability from BL Lacertae during the extreme flaring state of 2022

### 25 Mar 2026 v2

@@ 7 Conclusions
BL Lac has shown one of its three brightest -ray and multi-wavelength flares ever between September and November 2023. We have characterised its HE and VHE emission, extending up to a few TeV during the brightest periods of the flare. One of the key features of this event is the significant remarkably fast VHE -ray flux variability, with doubling flux timescales as short as 8.3 minutes, which sets tight constraints to the size of the VHE emitting region. This is also among the fastest flux variations ever detected at VHEs from blazars.
The HE-VHE spectrum shows a shift of the peak towards higher energies at the highest stages of the flare. Accompanying this behaviour we also report a softer-when-brighter evolution of the X-ray spectrum due to a clear transition of the X-ray emission from IC-dominated to synchrotron-dominated, meaning that the source also transitions from LSP to ISP, as seen in other periods for BL Lac in the past (Nilsson et al., 2018).
The broadband emission was modelled within a leptonic scenario, motivated by recent results reported by IXPE (Middei et al., 2023; Peirson et al., 2023; Agudo et al., 2025; Liodakis et al., 2025). One-zone models were not successful when trying to explain the broadband emission. Therefore, a co-spatial two-zone SSC model incorporating also the dusty torus EC contribution was used. Within this scenario, the evolution of the jet is driven by changes on the electron population of both regions, with injection of freshly accelerated electrons being the cause of the brightest outbursts. The minute-scale variability observed constrains the size of the compact blob region to be cm for the adopted Doppler factor , consistent with the value used here. This blob region is therefore the responsible of the intranight -ray variability, that can be associated with a plasmoid under a magnetic reconnection event.

### 07 Apr 2026 v3

@@ 7 Conclusions
BL Lac exhibited one of its three brightest -ray and multi-wavelength flares on record between September and November 2022. We characterised its HE and VHE emission, which extended up to a few TeV during the flares’ brightest periods. A key feature of this event is the remarkably fast VHE -ray flux variability, with flux doubling timescales as short as 8.3 min, which sets tight constraints on the size of the VHE emitting region. This represents one of the fastest flux variations ever detected at VHEs from blazars.
The HE-VHE spectrum shows a shift in the peak towards higher energies during the most intense stages of the flare. Accompanying this behaviour, we report a softer-when-brighter evolution of the X-ray spectrum resulting from a clear transition of the X-ray emission from IC-dominated to synchrotron-dominated. Consequently, the source transitions from LSP to ISP, as previously observed in BL Lac (Nilsson et al., 2018).
We modelled broadband emission within a leptonic scenario, motivated by recent IXPE results (Middei et al., 2023; Peirson et al., 2023; Agudo et al., 2025; Liodakis et al., 2025). One-zone models were unsuccessful in explaining the broadband emission. Therefore, we employed a co-spatial two-zone SSC model, also incorporating the EC contribution from the dusty torus. Within this scenario, the jet evolution is driven by changes in the electron populations of both regions, with injections of freshly accelerated electrons triggering the brightest outbursts. The observed minute-scale variability constrains the size of the compact blob region to cm for the adopted Doppler factor, , consistent with the value used in our model. This blob region is therefore responsible for the intranight -ray variability, which can be associated with plasmoids forming during a magnetic reconnection event.